Suppose an inflationary economy can be described by the following equations representing the goods and money markets: C = 20 + 0.7Yd M = 0.4Yd I = 70 – 0.1r T = 0.1Y G = 100 X = 20 Ld = 389 + 0.7Y – 0.6r Ls = 145 where G represents government expenditure, M is imports, X is exports, Y is national income, Yd is disposable income, T is government taxes (net of transfer payments), I is investment, r is the rate of interest, C is consumption, Ld is money demand, and Ls is money supply. (15) i) Use the inverse matrix method to solve for the equilibrium level of national income and the equilibrium rate of interest in this economy. (Note: ½ of the marks in this part are given for the correct set up of the equations. Explain what you are doing, including how equilibrium is established in each market.) (5) ii) Now use Cramer’s rule to find your answer
Answer #1

Goods market y= ct Itat x-m y = 20+ 0.7yd + 70 - 0.18 +100 + 20 - 0.494 Y = 20+ 0.37 - 0.18 Y = 210 + 0:3(4-0) -0.13 Y = 210IAL = T 0.73 10.70 on -0.6 = -0.438 -0.07 0.508 Ã = adja SEAL adiA = | cll (120 L cal (22 J Cij = (-)tj mij (11= (-nt (-0.621-0.6 -0.5081 -0.70 -ou 0.73 x= ÃB Pro x = - 1 2)0 1-0.6 L-0.70 – oil 0.7 / 2107 0.73 -zuu 0508 * 508 -oize T-126 + 24.4 -147((1 و ه علا = ۱A 0 . 43 0 . 1 ا کہ : ۵۰۰ | = - ۵۰ 558 دارم۱ و (2 | ۰۱ة | ہا ہا2 - | ای ۵۰ ۔ و .ہ 2 + 126 - = = ۱۵۰ - | ) 2 =